# MemoryFormer: Minimize Transformer Computation by Removing Fully-Connected Layers

### v1

@@ 3.4 Visualization
In a Memory Layer, we do not want most of the sub-vector $\mathbf{z}_{k}$ hashed to a few popular buckets while the rest of buckets in the table are rarely selected. We expect $\mathbf{z}_{k}$ to be hashed to all the buckets with the same probability. If so, the output space of the Memory Layer would be diversified enough to enlarge the capacity of MemoryFormer. We use 2048 sequences of 1024 token length to visualize the distribution of the frequency that each bucket is retrieved within a hash table in Figure 4. Specially, we choose the first table $\mathbf{T}_{1}$ and the last table $\mathbf{T}_{64}$ of the Q, K, V projection layers and the two layers in the FFN module from the first building block of the MemoryFormer-tiny model. As shown in Figure 4, the number of times each bucket is hashed to by $\mathbf{z}$ is generally uniform.

### v2

@@ 3.4 Visualization
In a Memory Layer, we do not want most of the sub-vector $\mathbf{z}_{k}$ hashed to a few popular buckets while the rest of buckets in the table are rarely selected. We expect $\mathbf{z}_{k}$ to be hashed to all the buckets with the same probability. If so, the output space of the Memory Layer would be diversified enough to enlarge the capacity of MemoryFormer. We use 2048 sequences of 1024 token length to visualize the distribution of the frequency that each bucket is retrieved within a hash table in Figure˜4. Specially, we choose the first table $\mathbf{T}_{1}$ and the last table $\mathbf{T}_{64}$ of the Q, K, V projection layers and the two layers in the FFN module from the first building block of the MemoryFormer-tiny model. As shown in Figure˜4, the number of times each bucket is hashed to by $\mathbf{z}$ is generally uniform.